OpenStudy (anonymous):

horizontal asymptote of: f(x) = (x^2 - 9) / (x - 2) horizontal asymptote of: f(x) = (x^2 - 9) / (x - 2) @Mathematics

OpenStudy (alfie):

\[\lim_{x->+\infty}\frac{x^2-9}{x-2} \rightarrow \lim_{x->+\infty}\frac{x^2}{x} = \infty \] So there is no horizontal asymptote.

OpenStudy (anonymous):

Oh wow, thats a lot of signs haha..

OpenStudy (anonymous):

there's a horizontal asymptote at x=2

OpenStudy (alfie):

That's a vertical one.

OpenStudy (anonymous):

oops.. good call haha... :P